OpenStudy (anonymous):

1260 inches squared=? feet squared Can you show me how to solve this

OpenStudy (johnweldon1993):

1260in^2 = ?ft^2 Well we know there are 12 inches in 1 foot...so we need to convert 1260 inches to feet... \[\large 1260\text{inches} \times \frac{1 \text{foot}}{12 \text{inches}} = \frac{1260\cancel{\text{inches}}\text{foot}}{12 \cancel{\text{inches}}} = 15 \text{feet}\]

OpenStudy (johnweldon1993):

Make sense?

OpenStudy (anonymous):

yes, 1260 inches squared = 15 feet squared

OpenStudy (johnweldon1993):

Perfect!

OpenStudy (anonymous):

Thanks for your help.